# Safe Explicable Policy Search

### v1

@@ Abstract
When users work with AI agents, they form conscious or subconscious expectations of them. Meeting user expectations is crucial for such agents to engage in successful interactions and teaming. However, users may form expectations of an agent that differ from the agent’s planned behaviors. These differences lead to the consideration of two separate decision models in the planning process to generate explicable behaviors. However, little has been done to incorporate safety considerations, especially in a learning setting. We present Safe Explicable Policy Search (SEPS), which aims to provide a learning approach to explicable behavior generation while minimizing the safety risk, both during and after learning. We formulate SEPS as a constrained optimization problem where the agent aims to maximize an explicability score subject to constraints on safety and a suboptimality criterion based on the agent’s model. SEPS innovatively combines the capabilities of Constrained Policy Optimization and Explicable Policy Search. We evaluate SEPS in safety-gym environments and with a physical robot experiment to show that it can learn explicable behaviors that adhere to the agent’s safety requirements and are efficient. Results show that SEPS can generate safe and explicable behaviors while ensuring a desired level of performance w.r.t. the agent’s objective, and has real-world relevance in human-AI teaming.
@@ I Introduction
Generally, users may not always be experts in AI systems, which may lead to them forming expectations that are different from the system’s planned behaviors. These differences may be due to, for instance, different understandings of the task objective and/or domain dynamics during decision-making. Explicable planning [2, 3] addresses the planning challenges where the AI agent aims to generate more expected plans or policies under such model differences. However, when the agent focuses on meeting user expectations, it may lead to safety issues.
A previous study [4], introduced the problem of Safe Explicable Planning (SEP) that aims to find user expected policies that are safe by considering a safety bound (provided by an expert) in the agent’s decision making (Fig. 1(b)). This work is limited to discrete states and actions, and to a planning setting where the models of the agent and the user (for generating expectations of the agent) are known. In reality, it is desirable for such methods to work in environments with continuous states and actions, and, more importantly, when the user’s model is not directly available. Our work bridges the gap by proposing a method for finding safe and explicable policies in continuous stochastic environments in a reinforcement learning (RL) setting. It is worth noting that, in a planning setting, the solution is a final plan or policy that is safe. However, in a learning setting, the solution may require the agent to be safe even during the learning process.
@@
Consider a motivating example, shown in Fig. 1(a), where an autonomous driving car is to merge with highway traffic. The safety criterion is to maintain a minimum distance from both the car immediately in front and the car immediately behind. The AI driving agent aims at saving energy (by reducing acceleration and deceleration) while the human values travel time more. The expected behavior is to merge as fast as it can. Even when assuming safety is assured, it may result in sudden and harsh acceleration and/or deceleration when the front and/or rear cars are very close to the AI agent. The optimal behavior to the agent is to slow down to wait for a better opportunity (where more distance in between the AI agent and the other cars are observed) to smoothly accelerate and merge. The safe explicable behavior may involve alternating between deceleration and acceleration for an opportunity (with moderate distances between the AI agent and other cars) to merge that does not require too much waiting. The proposed Safe Explicable Policy Search (SEPS) aims to best match the user’s expectation within the agent’s safety and quality limits, thus avoiding unsafe and inefficient behaviors during and after learning.
@@
Both safe behavior and explicable behavior generation, considered independently, can be addressed by the existing methods. For example, Constrained Markov Decision Process (CMDP) [5] and its solution techniques [6, 7, 8], etc., are effective in optimizing a task objective subject to safety constraints. The challenge in our problem is that the objective is defined over two different models with potentially different objectives and domain dynamics. This limits us from directly applying CMDP solutions. Similarly, a previous study on Explicable Policy Search (EPS) [3] proposes learning a surrogate reward function from user feedback on the agent’s behavior to inform learning. It is more practical than learning both the objective and domain dynamics and hence can be applied when the user’s model is unknown. However, EPS depends on the specification of a reconciliation parameter, and a weighting parameter for a few entropy terms, to combine an explicability score with the agent’s objective. It requires careful fine-tuning via trial and error due to the difficulty in relating these hyper-parameters to the importance of the objectives reconciled since they are at different scales (i.e., reward vs. entropy). We propose SEPS to combine the capabilities of CMDP (CPO [6]) and EPS while avoiding these limitations in addressing safe explicable behavior generation.
@@
Towards this end, we first formulate our objective based on EPS since it allows us to convert the consideration of two models to the consideration of two linearly combined objectives. This is particularly beneficial since it enables us to focus policy search on a single model, thus allowing us to apply CPO to the new objective. Meanwhile, to relax the requirement of the hyper-parameters, we adapt and reformulate the objective by breaking it into an objective and a constraint, where the constraint can be absolved into the CPO formulation, leading to a novel integration of CPO and EPS while avoiding their limitations for safe explicable behavior generation. The remaining challenge lies in the derivation of an analytic solution for the resulting constrained optimization problem, for which we provide a solution for the most common case with two constraints. Furthermore, we evaluate SEPS in different safety-gym environments and with a physical robot experiment. The results show that our approach is effective in searching for an explicable policy within the safety limits. Our approach also demonstrates its applicability to real-life scenarios in the robot experiment.
@@ II Related Work
With recent developments in AI, AI agents become accessible to various users where they either work alongside or interact with them. This has motivated advancements in Explainable AI (XAI) with the aim of building AI agents whose behavior is explainable to users [9, 10]. Our work is related to research that focuses on generating user-understandable behavior which is studied under different terminologies, viz., legibility [11], predictability [12], transparency [13], explicability [2], etc. For a detailed review of these studies, refer to [14].
@@
In the spectrum of Explainable AI literature, our work is most related to explicable planning [2, 3, 15, 16], a framework for finding a user expected behavior while balancing the costs incurred to the agent. These studies propose different methods to measure similarity to user expectations but do not consider safety issues. Other studies that consider human feedback in the agent’s learning process, such as  [17, 18], are capable of learning user preferred behavior in the presence of an expert designed RL objective. However, these methods assume the user’s belief of the domain dynamics is the same as that of the agent, which can lead to learning undesired behaviors [19]. Furthermore, these approaches do not consider safety issues either.
@@
Safe Explicable Planning (SEP) extends explicable planning to support the specification of a safety bound [4]. However, this work is limited to a planning setting where the decision models are given. In reality, such models may be difficult to provide a priori. Furthermore, it requires searching through the policy space, which is intractable in large, continuous MDPs. In this work, we consider the problem of finding a safe explicable policy in a reinforcement learning setting. However, we assume that the information regarding the user’s model is learned from user feedback and is given. This allows us to focus on developing the methodology for SEPS. Furthermore, unlike SEP, in our formulation, we consider the general case of continuous state and action spaces.
@@
In our work, we define the agent’s decision model as a Constrained Markov Decision Process (CMDP)  [5], which allows us to specify the costs associated with safety separately from the agent’s task objective. CMDPs are popularly used in Safe RL (refer to [20] for a review of Safe RL methods) due to the ease of safety specification. Several techniques have been proposed to solve safe RL formulated as CMDPs [6, 21, 22, 23]. However, these methods cannot be directly applied to solve SEPS due to its problem setting (refer to Fig. 1(b)), where the optimization is under two different models, specifically different domain dynamics models (or transition models in MDPs).
@@ III Preliminaries
A Markov Decision Process (MDP) is represented by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho,\gamma\rangle$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition probability function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is the reward function, $\rho(s)$ is the initial state distribution, and $\gamma$ is the discount factor. A stationary policy $\pi:\mathcal{S}\mapsto\mathcal{P}(\mathcal{A})$ is a map from states to probability distribution over actions, with $\pi(a|s)$ denoting the probability of selecting action $a$ in state $s$. We denote the set of all stationary policies by $\Pi$. The performance (expected discounted return) of a policy w.r.t. a measure $\mathcal{R}$ under the transition dynamics $\mathcal{T}$ is typically given by $J_{\mathcal{R},\mathcal{T}}(\pi)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t},s_{t+1})]$, where $\tau\sim\langle\pi,\mathcal{T}\rangle$ indicates that the distribution over trajectories depends on $\pi$ and $\mathcal{T}$ i.e., $s_{0}\sim\rho,a_{t}\sim\pi(.|s_{t}),s_{t+1}\sim\mathcal{T}(.|s_{t},a_{t})$. The probability of realizing a trajectory $\tau$ with policy $\pi$ is $p(\tau){=}\rho(s_{0})\prod_{t}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})$. Reinforcement learning algorithms (refer to [24]) seek to learn a policy that maximize the performance, i.e., $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi}J_{\mathcal{R},\mathcal{T}}(\pi)$.
@@
A constrained Markov Decision Process (CMDP) is an MDP augmented with constraints that restrict the permissible policies for the MDP. Specifically, $\mathcal{M}$ is augmented by auxiliary cost functions $\{\mathcal{C}_{i}\}_{i=1:k}$ and limits $\{d_{i}\}_{i=1:k}$ where each $\mathcal{C}_{i}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is a function, same as $\mathcal{R}$. The performance of a policy w.r.t. $\mathcal{C}_{i}$ is $J_{\mathcal{C}_{i},\mathcal{T}}(\pi)=E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}_{i}(s_{t},a_{t},s_{t+1})]$. Then, the set of all feasible stationary policies is denoted by $\Pi_{\mathcal{C}}\dot{=}\{\pi{\in}\Pi:\forall{i}J_{\mathcal{C}_{i},\mathcal{T}}(\pi)\leq d_{i}$}. The objective of a CMDP is to learn a feasible policy that maximizes the performance $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{\mathcal{C}}}J_{\mathcal{R},\mathcal{T}}(\pi)$.
@@
In explicable policy search (EPS) [3], two MDP models, the agent’s model $M_{A}$ and the user’s belief of the agent’s model $M_{A}^{H}$, are considered at play. The two MDPs share the same $\mathcal{S}$, $\mathcal{A}$, $\rho$, and $\gamma$ but may differ in their domain dynamics ($\mathcal{T}_{A}$, $\mathcal{T}_{A}^{H}$) and their reward functions ($\mathcal{R}_{A}$, $\mathcal{R}_{A}^{H}$). In EPS, the aim is to search for a policy that maximizes two objectives: the expected cumulative reward, and a policy explicability score, linearly weighted by a reconciliation factor ($\lambda$) i.e.,
@@
The explicability score (.$\mathcal{E}(.){=}-\mathcal{D}_{KL}(p_{A}(\tau)||p_{A}^{H}(\tau))$ is well motivated (refer to  [3]) and is defined as the negative KL-divergence of the trajectory distributions in the agent’s model and the human’s model denoted by $p_{A}(\tau)$ and $p_{A}^{H}(\tau)$, respectively. Essentially, the agent that maximizes $\mathcal{E}(.)$ would learn to replicate (as closely as possible) the human’s expected behavior of the agent.
Estimating the parameters in $p_{A}^{H}(\tau)$ (i.e. $\mathcal{T}_{A}^{H}$ and $\pi_{A}^{H}$), individually, is expensive. Instead, the authors propose to learn a surrogate reward function $u_{H}$ via a preference-based learning framework [25, 26] by presenting pairs of trajectories and asking human subjects which trajectory is expected. The learned reward function $u_{H}$ correlates with the distribution of human expected trajectories, $p_{A}^{H}(\tau)$. Furthermore, assuming a parametrized agent policy $\pi_{\theta}$, the surrogate reward signal and other log terms in the explicability measure effectively reshape the reward function in Eq. 1 and is approximated by the objective
$\theta^{*}\dot{=}\operatorname{arg\,max}_{\pi_{\theta}}\mathbb{E}_{p_{A}}\biggl{[}\sum_{t}\gamma^{t}(\mathcal{R}_{A}(s_{t},a_{t})\\ +\lambda(u_{H}(s_{t},a_{t})+\mathcal{H}_{\mathcal{T}_{A}}(s_{t+1}|s_{t},a_{t})+\mathcal{H}_{\pi_{\theta}}(a_{t}|s_{t})))\biggr{]},$ (2)
where $u_{H}(s_{t},a_{t})=\mathcal{T}_{A}^{H}(s_{t+1}|s_{t},a_{t})+\pi_{A}^{H}(a_{t}|s_{t})+C_{1}$, $\mathcal{H}_{\pi_{\theta}}$ is the policy entropy term (same as MaxEntRL [27]) which incentivizes the agent to explore uncertain actions, and $\mathcal{H}_{\mathcal{T}_{A}}$ is the environment entropy term which incentivizes the agent to explore parts of the environment that has more stochasticity.

### v3

@@ Abstract
When users work with AI agents, they form conscious or subconscious expectations of them. Meeting user expectations is crucial for such agents to engage in successful interactions and teaming. However, users may form expectations of an agent that differ from the agent’s planned behaviors. These differences lead to the consideration of two separate decision models in the planning process to generate explicable behaviors. However, little has been done to incorporate safety considerations, especially in a learning setting. We present Safe Explicable Policy Search (SEPS), which aims to provide a learning approach to explicable behavior generation while minimizing the safety risk, both during and after learning. We formulate SEPS as a constrained optimization problem where the agent aims to maximize an explicability score subject to constraints on safety and a suboptimality criterion based on the agent’s model. SEPS innovatively combines the capabilities of Constrained Policy Optimization and Explicable Policy Search to introduce the capability of generating safe explicable behaviors to domains with continuous state and action spaces, which is critical for robotic applications. We evaluate SEPS in safety-gym environments and with a physical robot experiment to show its efficacy and relevance in human-AI teaming.
@@ I Introduction
More specifically, users may not always be experts in AI systems, which may lead to them forming expectations that are different from the system’s planned behaviors. These differences may be due to, for instance, different understandings of the task objective and/or domain dynamics during decision-making. Explicable planning [2, 3] addresses the planning challenges where the AI agent aims to generate more expected plans or policies under such model differences. However, when the agent focuses on meeting user expectations, it may lead to safety and/or efficiency issues.
@@
Consider a motivating example, shown in Fig. 1(a), where an autonomous driving car is to merge with highway traffic. The autonomous driving agent seeks to conserve energy so that it avoids hard acceleration and deceleration. However, such an objective can directly conflict with the user’s expectation of minimizing the travel time. More specifically, while the optimal behavior of the agent is to slow down and wait for a large opening between the front and rear cars, the expected behavior is to merge as fast as it can, which can lead to safety risks. We can incorporate a safety criterion to maintain a minimum distance from the other cars to guarantee safety. However, an agent seeking explicable behaviors subject to the safety constraint may still opt for hard acceleration and deceleration with relatively small openings. A safe explicable behavior may involve alternating between soft deceleration and acceleration for an opportunity with a larger opening. The proposed Safe Explicable Policy Search (SEPS) aims to best match the user’s expectation within the agent’s safety and quality limits, thus avoiding unsafe and inefficient behaviors during and after learning.
A previous study [4], introduced the problem of Safe Explicable Planning (SEP) that aims to find user expected policies that are safe by considering a safety bound (provided by an expert) in the agent’s decision making (Fig. 1(b)). This work is limited to discrete states and actions, and to a planning setting where the models of the agent and the user (for generating expectations of the agent) are known. In reality, it is desirable for such methods to work in environments with continuous states and actions, and, more importantly, when the user’s model is not directly available. Our work bridges the gap by proposing a method for finding safe and explicable policies in continuous stochastic environments in a reinforcement learning (RL) setting, which is especially valuable for robotic applications. It is worth noting that, in a planning setting, the solution is a final plan or policy that is safe. However, in a learning setting, the solution may require the agent to be safe even during the learning process.
@@
Both safe behavior and explicable behavior generation, considered independently, can be addressed by the existing methods. For example, Constrained Markov Decision Process (CMDP) [5] and its solution techniques [6, 7, 8], etc., are effective in optimizing a task objective subject to safety constraints. The challenge in our problem is that the objective is defined over two different models with potentially different objectives and domain dynamics. This limits directly applying CMDP solutions. Similarly, a previous study on Explicable Policy Search (EPS) [3] proposes learning a surrogate reward function from user feedback on the agent’s behavior to inform learning. It is more practical than learning both the reward and domain dynamics and hence can be applied when the user’s model is unknown. However, EPS depends on the specification of a reconciliation parameter, and a weighting parameter for a few entropy terms, to combine an explicability score with the agent’s objective. It requires careful fine-tuning via trial and error due to the difficulty in relating these hyper-parameters to the importance of the objectives reconciled since they are at different scales (i.e., reward vs. entropy). These issues in existing methods limit their applicability to the generation of safe and explicable behaviors.
@@
To take advantage of both sides for safe explicable behavior generation, we first formulate our objective based on EPS since it allows us to convert the consideration of two models to the consideration of two linearly combined objectives. This is particularly beneficial since it enables us to focus policy search on a single model, reducing the new objective to a CMDP. Meanwhile, to relax the requirement of the hyper-parameters, we adapt and reformulate the objective by breaking it into an objective and a constraint, where the constraint can be dissolved into the CMDP formulation. We consider the popular Constrained Policy Optimization (CPO) [6] approach to solve the reduced CMDP problem. Our work considers a novel integration of EPS and CPO, while avoiding their limitations, to introduce safe explicable policy search. The remaining challenge lies in the derivation of an analytic solution for the resulting constrained optimization problem, for which we provide a solution for the most common case with two constraints. Furthermore, we evaluate SEPS in different safety-gym environments and with a physical robot experiment. The results show that our approach is effective in searching for an explicable policy within the safety limits. Our approach also demonstrates its applicability to real-life scenarios in the robot experiment.
@@ II Related Work
With recent developments in AI, AI agents become accessible to various users, where they either work alongside or interact with them. This has motivated advancements in Explainable AI (XAI) with the aim of building AI agents whose behavior is explainable to users [9, 10]. Our work is related to research that focuses on generating user-understandable behavior which is studied under different terminologies, viz., legibility [11], predictability [12], transparency [13], explicability [2], etc. For a detailed review of these studies, refer to [14].
@@
In the spectrum of Explainable AI literature, our work is most related to explicable planning [2, 3, 15, 16], a framework for finding a user expected behavior while balancing the costs incurred to the agent. These studies propose different methods to measure similarity to user expectations but do not consider safety issues. Other studies that consider human feedback in the agent’s learning process, such as  [17, 18, 19], are capable of learning user-preferred behavior in the presence of an expert designed RL objective. However, these methods assume the user’s belief of the domain dynamics is the same as that of the agent, which can lead to learning undesired behaviors [20] when untrue. Furthermore, these approaches do not consider safety issues either.
@@
Safe Explicable Planning (SEP) extends explicable planning to support the specification of a safety bound [4]. However, it requires searching through the policy space, which is intractable in large, continuous MDPs. Furthermore, SEP is limited to a planning setting in which the decision models are given. In reality, such models may be difficult to provide a priori. The models can be learned individually, for example, [21] learned the domain dynamics and [22] learned the reward function, or jointly  [23, 20]. However, for real-world problems, this can be challenging due to its complexity and high dimensionality. In this work, we consider the problem of finding a safe, explicable policy in a reinforcement learning setting. A prior work [3] learns sufficient information about the user’s model from user feedback, through a surrogate reward model. In this work, we assume that the surrogate reward model is given. This allows us to focus on developing the methodology for SEPS. Furthermore, unlike SEP, in our formulation, we consider the general case of continuous state and action spaces.
@@
In our work, we define the agent’s decision model as a Constrained Markov Decision Process (CMDP)  [5], which allows us to specify the costs associated with safety separately from the agent’s task objective. CMDP is a popularly used formulation in Safe RL (refer to [24] for a review of Safe RL methods). Another common way to handle safety in RL is to adopt a risk-aware formulation that includes a risk measure in the optimization criterion, which may take various forms, such as the exponential utility function  [25, 26], a weighted sum of the return and risk functions  [27, 28], and probabilistic conditions  [29, 30]. In this work, we adopt the constrained formulation due to the ease of specification and its safety guarantees. Several techniques have been proposed to solve a CMDP [6, 31, 32, 33]. However, these methods cannot be directly applied to solve SEPS due to its problem setting (refer to Fig. 1(b)), where the optimization is under two different MDPs, specifically different transition models in MDPs. We consider the Constrained Policy Optimization (CPO) [6] approach to solve CMDP due to its convergence and safety guarantees. Other recent studies  [34, 35, 36] that build on top of CPO can be easily applied to solve SEPS.
@@ III Preliminaries
Our formulation is based on Markov Decision Processes (MDP). An MDP is represented by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho,\gamma\rangle$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition probability function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is the reward function, $\rho:\Delta(\mathcal{S)}$ is the initial state distribution, and $\gamma$ is the discount factor. A stationary policy $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ is a map from states to probability distribution over actions, with $\pi(a|s)$ denoting the probability of selecting action $a$ in state $s$. We denote the set of all stationary policies by $\Pi$. The performance (expected discounted return) of a policy w.r.t. a measure $\mathcal{R}$ under the transition dynamics $\mathcal{T}$ is typically given by $J_{\mathcal{R},\mathcal{T}}(\pi)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t},s_{t+1})]$, where $\tau{\sim}\langle\pi,\mathcal{T}\rangle$ indicates that the distribution over trajectories depends on $\pi$ and $\mathcal{T}$ i.e., $s_{0}{\sim}\rho,a_{t}{\sim}\pi(.|s_{t}),s_{t+1}{\sim}\mathcal{T}(.|s_{t},a_{t})$. The probability of realizing a trajectory $\tau$ with policy $\pi$ is $p(\tau){=}\rho(s_{0})\prod_{t}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})$. Reinforcement learning algorithms (refer to [37]) seek to learn a policy that maximize the performance, i.e., $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi}J_{\mathcal{R},\mathcal{T}}(\pi)$.
@@
CMDP and Constrained Policy Optimization (CPO): A CMDP is an MDP augmented with constraints that restrict the permissible policies for the MDP. Specifically, $\mathcal{M}$ is augmented by auxiliary cost functions $\{\mathcal{C}_{i}\}_{i=1:k}$ and limits $\{d_{i}\}_{i=1:k}$ where each $\mathcal{C}_{i}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is a function, same as $\mathcal{R}$. The performance of a policy w.r.t. $\mathcal{C}_{i}$ is $J_{\mathcal{C}_{i},\mathcal{T}}(\pi)=E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}_{i}(s_{t},a_{t},s_{t+1})]$. Then, the set of all feasible stationary policies is denoted by $\Pi_{\mathcal{C}}\dot{=}\{\pi{\in}\Pi:\forall{i}J_{\mathcal{C}_{i},\mathcal{T}}(\pi)\leq d_{i}$}. The objective of a CMDP is to learn a feasible policy that maximizes the performance $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{\mathcal{C}}}J_{\mathcal{R},\mathcal{T}}(\pi)$.
CPO [6] is a popular approach to solving CMDP. CPO applies trust region policy optimization (TRPO) [38] to policy search updates in a CMDP. The policy update rule is given by,
$\pi^{n+1}=\operatorname{arg\,max}_{\pi\in\Pi_{\theta}}\mathbb{E}_{\pi^{n},\mathcal{T}}[A_{\mathcal{R}}^{\pi^{n}}(s,a)]\\ s.t.J_{\mathcal{C}_{i},\mathcal{T}}(\pi_{n})+\frac{1}{1-\gamma}\mathbb{E}_{\pi^{n},\mathcal{T}}[A_{\mathcal{C}_{i}}^{\pi^{n}}(s,a)]\leq d_{i}\hskip 10.0pti=1,...,k\\ \bar{\mathcal{D}}_{KL}(\pi||\pi_{n})\leq\delta\hskip 60.0pt$ (1)
where, $A_{.}^{\pi}(s,a)\dot{=}Q_{.}^{\pi}(s,a)-V_{.}^{\pi}(s)$ is the advantage function given by the on-policy value function $V_{.}^{\pi}(s)$ and the on-policy action-value function $Q_{.}^{\pi}(s,a)$, $\Pi_{\theta}\subset\Pi$ is a set of parameterized policies with parameters $\theta$, $\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})=\mathbb{E}_{s\sim\pi^{n}}[\mathcal{D}_{KL}(\pi||\pi_{n})[s]]$ is a measure of KL Divergence between the current policy $\pi_{n}$ and a potential new policy $\pi$, and $\delta>0$ is the step size. The set of policies that satisfy the third constraint is called the trust region, i.e. $\{\pi_{\theta}\in\Pi_{\theta}|\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})\leq\delta\}$. For performance guarantees and constraint satisfaction guarantees, see [38] and [6], respectively.
For continuous MDPs with high dimensional spaces, solving Eq. (1) exactly is intractable and requires an approximation. The CPO method proposes the following approximation
$\theta_{n+1}=\operatorname{arg\,max}_{\theta}g^{T}(\theta-\theta_{n})\\ s.t.\hskip 3.0ptc_{i}+b_{i}^{T}(\theta-\theta_{n})\leq 0\hskip 10.0pti=1,...,k\\ \frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta\hskip 80.0pt$ (2)
where, $g$ is the gradient of the objective (explicability measure), $b_{i}$ is the gradient of the constraint $i$, $c_{i}\dot{=}J_{\mathcal{C}_{i},\mathcal{T}}(\pi_{n})-d_{i}$, and $H$ is the Hessian of the KL Divergence.
@@
Explicable Policy Search (EPS): In EPS [3], two MDP models, the agent’s model $M_{A}$ and the user’s belief of the agent’s model $M_{A}^{H}$, are considered at play. The two MDPs share the same $\mathcal{S}$, $\mathcal{A}$, $\rho$, and $\gamma$ but may differ in their domain dynamics ($\mathcal{T}_{A}$, $\mathcal{T}_{A}^{H}$) and their reward functions ($\mathcal{R}_{A}$, $\mathcal{R}_{A}^{H}$). In EPS, the aim is to search for a policy that maximizes a linear combination of two objectives, namely, the expected cumulative reward and a policy explicability score, weighted by a reconciliation factor ($\lambda$):
@@
The explicability score $\mathcal{E}(.)$ is well motivated (refer to  [3]) and is defined as the negative KL-divergence of the trajectory distributions in the agent’s model and the human’s model denoted by $p_{A}(\tau)$ and $p_{A}^{H}(\tau)$, respectively. Formally, $\mathcal{E}(.)=-\mathcal{D}_{KL}(p_{A}(\tau)||p_{A}^{H}(\tau))$ $\mathbb{E}_{p_{A}}[\sum_{t}\log\mathcal{T}_{A}(s_{t+1}|s_{t},a_{t})+\log\pi_{A}(a_{t}|s_{t})-\log\mathcal{T}_{A}^{H}(s_{t+1}|s_{t},a_{t})-\log\pi_{A}^{H}(a_{t}|s_{t})]+C$. Essentially, the agent that maximizes $\mathcal{E}(.)$ would learn to replicate (as closely as possible) the human’s expected behavior of the agent.
Estimating the parameters in $p_{A}^{H}(\tau)$ (i.e. $\mathcal{T}_{A}^{H}$ and $\pi_{A}^{H}$), individually, is expensive. Instead, the authors propose to learn a surrogate reward function $u_{H}$ via a preference-based learning framework [39, 40] by presenting pairs of trajectories and asking human subjects which trajectory is expected. The goal of learning $u_{H}$ is to obtain feedback on the user’s assessment of explicableness of the agent’s behavior, which is evaluated under the user’s belief (i.e., $\mathcal{M}_{A}^{H}$). The learned reward function $u_{H}$ correlates with the distribution of human expected trajectories i.e., $u_{H}(s_{t},a_{t})=\log\mathcal{T}_{A}^{H}(s_{t+1}|s_{t},a_{t})+\log\pi_{A}^{H}(a_{t}|s_{t})+C_{1}$. Assuming a parametrized agent policy $\pi_{\theta}$, the surrogate reward and other log terms in the explicability score effectively reshape the reward function $\mathcal{R}_{A}$, and Eq. (3) is approximated as follows
$\theta^{*}\dot{=}\operatorname{arg\,max}_{\pi_{\theta}}\mathbb{E}_{p_{A}}\Bigl[\sum_{t}\gamma^{t}(\mathcal{R}_{A}(s_{t},a_{t})\\ +\lambda(u_{H}(s_{t},a_{t})+\mathcal{H}_{\mathcal{T}_{A}}(s_{t+1}|s_{t},a_{t})+\mathcal{H}_{\pi_{\theta}}(a_{t}|s_{t})))\Bigr],$ (4)
where $\mathcal{H}_{\pi_{\theta}}$ and $\mathcal{H}_{\mathcal{T}_{A}}$ are the entropies of the agent’s policy and domain dynamics (i.e., the true environment dynamics), respectively.